OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x - cos x

OpenStudy (anonymous):

Consider that: \[ \sin(3x) =\sin(x+2x) = \sin(x+(x+x)) \]And that: \[ \sin(x+y) = \sin(x)\cos(y)+\sin(y)\cos(x) \]

OpenStudy (anonymous):

Also:\[ \cos(x+y) = \cos(x)\cos(y) -\sin(x)\sin(y) \]

OpenStudy (anonymous):

I believe i have to rewrite it using Sin(2x) = 2sin(x)cos(x) This is what i have Sin(3x)-cos(x) Sin(x+2x) - cos(x) sin(x)+2sin(x)-cos(x) sin(x)+2sin(x)Cos(x)-cos(x)

OpenStudy (anonymous):

lost after that

OpenStudy (anonymous):

You made a mistake: \[ \sin(x+2x) - \cos(x) \neq \sin(x)+2\sin(x)-\cos(x) \]

OpenStudy (anonymous):

You have to use: \[ \sin(x+y) = \sin(x)\cos(y)+\sin(y)\cos(x) \]You should try letting \(y=2x\), and going from there.

OpenStudy (anonymous):

would it be Sin(x)cos(2x)+sin(2x)cos(x)-cos(x) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now you have to use it again.

OpenStudy (anonymous):

This equation: \[ \sin(2x) = 2\sin(x)\cos(x) \]Is really just letting \(y = x\) in this equation: \[ \sin(x+y) = \sin(x)\cos(y)+\sin(y)\cos(x) \]Since: \[ \begin{split} \sin(2x) = \sin(x+x) = \sin(x)\cos(x)+\sin(x)\cos(x) = 2\sin(x)\cos(x) \end{split} \]

OpenStudy (anonymous):

So then i Have Sin(x)-2sin^3(x) + sin(x)cos(x)+sin(x)cos(x)-cos(x)??

OpenStudy (anonymous):

You should be showing your work.

OpenStudy (anonymous):

sin(x)(1-2sin^2(x))+sin(x)cos(x)+sin(x)cos(x)-cos(x) is where i got it from, changed the cos(2x) to (1-2sin^2(x)) in the formula

OpenStudy (anonymous):

or would it be sin(x)cos(2x)+sin(2x)cos(x)-cos(x) sin(x)cos(x)cos(x)-sin(x)sin(x)+sin(x)cos(x)+sin(x)cos(x)cos(x)-cos(x)???

OpenStudy (anonymous):

these are my options 1) 2 sin x - sin^3(x) - cos x 2) 2 sin x cos^2(x) + sin x - 2 sin^3(x) - cos x 3) 2 sin^3(x) cos^4(x) + 1 4) 3 sin x cos x - sin^3(x) - cos x